| Operation Summary for February 10, 2012 - February 13, 2012 |
|
NuMI reported receiving 5.59E18 protons on target for the period from 2/6/12 to 2/13/12. Friday February 10 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam. During the day shift, at 8:23 AM, a Booster RF station (BRF18) tripped off on a LCW off indication; Operators investigated and reset the station. Linac personnel reported at 8:53 AM that they were working on a Klystron RF station (KRF2); beam returned at 10:16 AM. LRF3 tripped off at 10:16 AM; it reset. Switchyard (SY) personnel began a two-hour high-energy study to the Meson dump. A MI person began a 30-minute study at 11:15 AM. SY personnel began a Meson M01 split beam study at 12:56 PM. At 1:42 PM, a SY power supply (S:V204) tripped off on a high temperature indication; a power supply person investigated and had the supply back on within 30 minutes. At 1:46 PM, a Muon Rings study was interrupted due to bad beam positions; Operators had beam back for the study within ten minutes. During the evening shift, beam studies for the Meson MTest area ended at 4:16 PM. At 4:35 PM, BRF11 tripped off; it reset. SY personnel ended their split beam studies at 4:38 PM. The NuMI beam permit tripped off at 5:03 PM due to bad beam positions; it reset. At 5:56 PM, KRF1 tripped off; it reset. Operators conducted a search and secure of the SeaQuest Neutrino Area enclosure NM2 and reported finding a small LCW leak on a quadrupole magnet (NM2Q1-3) at 6:53 PM; Operators passed this on to the Water Group. At 8:04 PM, LRF4 tripped off for the first of two times in about 30 minutes; it reset. Operators completed a Booster injection tune at 8:42 PM. Saturday February 11 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 1:11 AM, a Linac quadrupole power supply (L:QPS104) tripped off; it reset. At 1:41 AM, MIRF10 tripped off; it reset. A Booster vacuum pump (IPQ16) tripped off; it reset. A MI kicker (I:KPS4E) tripped off at 4:37 AM on a pre-fire indication; it reset. LRF4 tripped off at 4:38 PM; it reset. At 4:42 AM, BRF4 tripped off; it reset. Operators reported at 6:43 AM that they had completed a 750 KeV line tune and a Booster injection tune. During the day shift, Operators had the Muon Rings ready for an access at 8:27 AM. At 12:09 PM, L:QPS104, 105, and 108 all tripped off; they reset. At 12:12 PM, MIRF16 tripped off; it reset. At 12:19 PM, LRF2 tripped off; it reset. Operators completed a Booster injection tune at 3:48 PM. During the evening shift, at 4:05 PM, S:V204 tripped off on a power supply high temperature indication; it reset. At 5:39 PM, KRF5 tripped off; it reset. At 6:32 PM, BRF6 tripped off; it reset. At 11:32 PM, LRF5 tripped off; it reset. Operators loaded the Recycler with protons at 11:36 PM. Sunday February 12 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 1:08 AM, LRF4 tripped off; it reset. At 1:12 AM, L:QPS108 tripped off; it reset. Operators reported at 3:28 AM that they had adjusted the low energy Buncher phase module and got a large increase in Booster efficiency. Operations completed a Booster injection tune at 4:24 AM. BRF7 tripped off at 7:27 AM; it reset. During the day shift, at 1:33 PM, BRF11 tripped off; it reset. During the evening shift, Operators reported at 4:19 PM that the MI South Anode Power Supply tripped off; Operators reset the supply but then MIRF2 tripped off on a control power fault indication. The station reset. KRF1 tripped off at 5 PM; it reset. MIRF16 tripped off at 5:09 PM; it reset. At 6:01 PM, all the NM2 power supplies tripped off with two of the power supplies showing a clock error indication; everything reset. LRF1 tripped off at 8:17 PM; it reset. At 9:11 PM, Operations prepared NM2 for an access on Monday morning. BRF11 tripped off at 10:20 PM; it reset. KRF5 tripped off at 11:24 PM; it reset. BRF14 tripped off at 11:26 PM; it reset. Monday February 13 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, Operations completed at 750 KeV line tune at 12:16 AM. LRF4 tripped off at 4:21 AM; it reset. At 5:32 AM, Operators reported a LCW leak and began a search of the Booster Galleries, the Transfer Hall Gallery, Linac upper and lower levels, the SY utility tunnel, the area leading to CUB underneath the Booster, and the PA test room, but found nothing. They checked the SY service building and the Neutrino Area G2 enclosure. After talking with a Muon Rings person, Operators checked the service buildings and then, at 7:01 AM, reported finding the leak in the AP30 tunnel on a magnet (A3B3); they valved the magnet LCW off and called in Water Group personnel. Operations Coordinator (OpsCo) The OpsCo agrees with the Booster and MI personnel plans to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 02/13/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |